If 𝑭 = πŸ‘π’™π’šπ’Š βˆ’ π’š 𝟐 𝒋 evaluate ∫ 𝑭 βˆ™ 𝒅𝒓 where C is the curve in the xy plane, π’š = πŸπ’™ 𝟐 , from (0,0) to (1,2)


1
Expert's answer
2021-06-02T11:54:43-0400

Given "\vec F=3xy\vec i-y^2\vec j"

"\vec r=x\vec i+y\vec j=>\vec {dr}=dx\vec i+dy\vec j"


"F\cdot \vec dr=3xydx-y^2dy"

"C" is the part of the curve "y=2x^2" from"(0, 0)" to "(1, 2)."


"x(t)=t, y(t)=2t^2, t\in[0, 1]"

"dx=dt, dy=4tdt"


Then


"\int_{C}\vec F\cdot \vec dr=\displaystyle\int_{0}^{1}(3t(2t^2)-(2t^2)^2(4t))dt"

"=\displaystyle\int_{0}^{1}(6t^3-16t^5)dt=\big[\dfrac{3t^4}{2}-\dfrac{8t^6}{3}\big]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=-\dfrac{7}{6}"